OpenStudy (anonymous):

What is the solution of the following system? -3x-2y=-12 9x+6y=-9

OpenStudy (cwrw238):

multiply first equation by 3 and add what do u get?

OpenStudy (anonymous):

Huh?

OpenStudy (cwrw238):

-3x - 2y = -12 * by 3:- -9x - 6y = -36 now add this to 9x + 6y = -9

OpenStudy (anonymous):

k so how do i find the solution here....?

OpenStudy (cwrw238):

adding 2 equation gives 0 = -45

OpenStudy (jiteshmeghwal9):

don't mind on my replies concentrate on @cwrw238 's reply

OpenStudy (cwrw238):

which is absurd so there are no solutions to this system

OpenStudy (anonymous):

(2, 1) no solutions (–2, –1) infinitely many solutions

OpenStudy (jiteshmeghwal9):

this equation don't exist

OpenStudy (anonymous):

those are the multiple choice..

OpenStudy (anonymous):

Okkk got it.. Thanks!

OpenStudy (cwrw238):

yw